# X-ray bright AGN in local dwarf galaxies: insights from eROSITA

### v1

@@ Abstract
Although supermassive black holes (SMBHs) reside in the heart of virtually every massive galaxy, it remains debated whether dwarf galaxies also commonly host SMBHs. Because low-mass galaxies may retain a memory of the assembly history of their black holes, probing the black hole occupation fraction of local dwarf galaxies might offer insights into the growth and seeding mechanisms of the first black holes. In this work, we exploit the Western half of the eROSITA all-sky survey (covering $20,000~{}\rm{deg^{2}}$) and compile a catalog of accreting SMBHs in local ($D<200$ Mpc) dwarf galaxies. After cleaning our sample from cosmic X-ray background sources, X-ray binaries, and ultraluminous X-ray sources, we identify 74 AGN-dwarf galaxy pairs. Using this large and uniform sample, we derive a luminosity function of dwarf galaxy AGN, fitting it with a power law function and obtaining ${\rm d}N/{\rm d}L_{\textup{X}}=(15.9\pm 2.2)\times L_{\textup{X}}^{-1.63\pm 0.05}$. Measuring the offset between the centroid of dwarf galaxies and the X-ray sources, we find that about $50\%$ of the AGN are likely off-nuclear, in agreement with theoretical predictions. We also compare the black hole-to-stellar mass relation of the AGN in our sample with the local and high-redshift relations, finding that our sources better adhere to the former. This suggests that local AGN across different mass scales underwent a similar growth history. Finally, we compare our sources with semi-analytical models: while our sample is too shallow to distinguish between different seeding models, it favors a growth mechanism linked to the star-formation rate of the host galaxy.
@@ 1 Introduction
Alternative methods to identify SMBHs in dwarf galaxies rely on optical, UV, and infrared variability (Baldassare et al., 2018, 2020; Martínez-Palomera et al., 2020; Ward et al., 2022; Arcodia et al., 2024), coronal lines (Satyapal et al., 2007; Cann et al., 2018, 2021; Salehirad et al., 2022; Reefe et al., 2023), radio emission (Mezcua et al., 2018; Reines et al., 2020; Davis et al., 2022), and X-ray observations (Schramm et al., 2013; Lemons et al., 2015; Pardo et al., 2016; Mezcua et al., 2018; Birchall et al., 2020; Latimer et al., 2021; Bykov et al., 2024). These last two methods are particularly effective in identifying AGN due to the high nuclear-to-host contrast that these bands offer: high angular resolution radio observations allow the separation of the AGN from other radio-emitting components of the galaxy, and the X-ray spectrum of AGN is distinct from other stellar X-ray sources (Ho, 2008; Merloni, 2016). However, a sensitive volume-limited radio survey of dwarf galaxies in the nearby universe is prohibitively expensive (Reines et al., 2020). Similarly, X-ray observational campaigns have, to date, been either too shallow and with poor angular resolution, or too limited in sky coverage.
The game-changer in this domain is eROSITA (extended ROentgen Survey with an Imaging Telescope Array, Predehl et al. 2021), the main instrument aboard the German-Russian satellite Spektrum Roentgen Gamma (SRG) (Sunyaev et al., 2021). eROSITA provides today’s most sensitive all-sky survey in the X-ray band, and eRASS1, the first data release, comprises more than 930,000 individual sources across the Western half of the sky (Merloni et al., 2024). In this work, we leverage the opportunity offered by eRASS1 to build a sample of uniformly selected AGN in local dwarf galaxies, thereby probing their black hole occupation fraction.
A similar effort has been conducted by Bykov et al. (2024), who utilized eROSITA data from the Eastern half of the sky and the MPA-JHU galaxy catalog, based on the SDSS DR12 (Alam et al., 2015), with promising results: they obtained a sample of 99 X-ray source-dwarf galaxy pairs, of which 82 were identified as AGN.
@@ 2.3 Unresolved X-ray binaries
The formula provided by Lehmer et al. (2010) derives the luminosity in the $2-10$ keV band, so we converted the $0.2-2.3$ keV luminosity provided by the eRASS1 catalog to the $2-10$ keV band assuming a Galactic absorption of $3\times 10^{20}$ cm-2 and a power law with photon index $\Gamma=1.8$.
@@ 2.4 Resolved ultra-luminous X-ray sources
Figure 5: A representative subset of our sample of AGN hosted in local dwarf galaxies. The three-color composite images were obtained using the SDSS gri-filters. The X-ray counts are shown with the white (two counts) and red (one count) pixels. The cross and the small magenta circle indicate the centroid and positional uncertainty ($\approx 5^{\prime\prime}$) of the X-ray source taken from the eRASS1 point source catalog (Merloni et al., 2024), respectively. The large cyan circle shows the PSF size ($r=26^{\prime\prime}$) of eROSITA in scanning mode.
@@ 2.6 Final sample of dwarf galaxy-X-ray source pairs
Our final sample, after removing the four known ULXs, is composed of 74 dwarf galaxies with an eROSITA-detected source. In Figure 6 we show the X-ray luminosity of the sources as a function of the redshift of the galaxies. Postcard images of a representative set of three of these candidates are shown in Figure 5. All the relevant properties of each dwarf galaxy-X-ray source pair are reported in Table 2.
@@ 3.2 Off-nuclear SMBHs
These results align with theoretical predictions and observational evidence (Bellovary et al., 2019; Reines et al., 2020; Bellovary et al., 2021) for AGN in dwarf galaxies (as also shown in blue in Figure 7), although some caveats must be considered. The large PSF of eROSITA in scanning mode ($\approx 26\arcsec$) prevents a precise location of the X-ray sources, and dwarf galaxies often exhibit irregular morphology, making it sometimes difficult to identify their centers (Pasetto et al., 2003; Lazar et al., 2024). To conclusively determine the location of the X-ray sources, follow-up high spatial resolution observations with Chandra would be needed.
@@ 3.4 Galaxy-to-black hole mass relation
Exploiting the analogies between high-redshift AGN and AGN hosted in low-mass galaxies, we further investigate these effects by studying the $M_{\bullet}-M_{\star}$ relation in our dwarf galaxy sample. We infer the SMBH mass from the X-ray luminosity of the AGN. To this end, we employ the bolometric correction $L_{\textup{bol}}/L_{\textup{2-10keV}}=16.7$ (see e.g., Lusso et al. 2012; Duras et al. 2020) and assume that all black holes in our sample accrete at $1\%$ of their Eddington ratio. The inferred black hole masses are in the range of $\sim 10^{4}-10^{1}0^{7}\ \rm{M_{\odot}}$. Figure 9 shows the $M_{\bullet}-M_{\star}$ relation for the AGN in our sample. The blue line and shaded region indicate the local relation for AGN and its dispersion (Reines & Volonteri, 2015). The red and green lines and shaded regions show the high-redshift relations proposed by Pacucci et al. (2023) and Li et al. (2024), respectively.
Our sample of AGN in dwarf galaxies is compatible with the low-redshift $M_{\bullet}-M_{\star}$ relation, although it shows significant scatter and is biased towards higher luminosities at the low-mass end of the plot (the dotted line indicates the lower limit on the detectable black hole mass). This suggests that local AGN hosted in dwarf galaxies share the same $M_{\bullet}-M_{\star}$ relation as more massive $z=0$ galaxies, in agreement with previous results (Schutte et al., 2019).
@@ 4 Comparison with semi-analytical models
Semi-analytical models (SAM) represent a powerful tool for studying BH seeding mechanisms as well as for exploring growth modes of BHs (Somerville et al., 2008; Bower et al., 2010; Ricarte & Natarajan, 2018a, b). In this work, we compare the observed BH occupation fraction with mock catalogs of AGN built following the approach described in Chadayammuri et al. (2023) based on the SAMs developed by Ricarte & Natarajan (2018b).
These SAMs consider two seeding mechanisms and three different mass growth modes, combined to build six mock catalogs. The seeding mechanisms comprise either rare “heavy” seeds, typically resulting from direct collapse (DC) scenarios, distributed following Lodato & Natarajan (2006) and Volonteri & Begelman (2010), or abundant “light” seeds with a distribution motivated by cosmological simulations (Hirano et al., 2015). This second scenario is motivated by cosmological simulations where BHs are produced by the death of population III (popIII) stars and the mass distribution of the seeds is represented by a power law spanning from 30 to 100 solar masses (Hirano et al., 2015).
Each seeding mechanism is coupled with one of three mass growth modes, detailed below.
@@
  • Broad Line Quasar (BLQ): When a major merger occurs, central BHs grow at Eddington ratios drawn from observed broad line quasars, allowing occasional super-Eddington accretion. Otherwise, the accretion rate is zero. [^2] This was shown to be adequate for reproducing quasar luminosity functions (Ricarte et al., 2019), but fails to produce enough faint AGN (Chadayammuri et al., 2023).
We can probe growth modes and seeding mechanisms by comparing our observed AGN properties with the mock catalogs built upon the described SAMs. Following Chadayammuri et al. (2023), for each SAM, we build the corresponding mock catalog by computing the expected X-ray luminosity based on the mass and mass accretion rate of each BH. We assume an efficiency in converting mass accretion rate into luminosity of $10\%$ and a bolometric-to-X-ray luminosity correction of $10\%$, which are both common and adequate choices (see e.g., Lusso et al. 2012; Duras et al. 2020). We then filter the obtained mock catalogs using the same luminosity, mass, and distance criteria adopted in the construction of our AGN sample in dwarf galaxies. This approach allows for a direct comparison between the filtered mock catalogs and our observed sample.
Figures 10 and 11 present our main results, showing the normalized distributions of host-galaxy masses and luminosities for our sample (in grey) and the six mock catalogs built on different models (colored dashed lines). Three models roughly reproduce the distribution of observed host-galaxy mass: DC+AGNms, popIII+AGNms, and popIII+BLQ. However, the popIII+BLQ model fails to predict the observed distribution of AGN luminosities. Even the two best-performing models (DC+AGNms and popIII+AGNms) fail to reproduce the observed luminosity distribution. They predict distributions heavily skewed toward high luminosities, peaking at about $10^{41.5}\ \rm{erg\ s^{-1}}$, while the observed distribution peaks roughly at $10^{39.5}\ \rm{erg\ s^{-1}}$. This suggests that the preferred accretion mode is the one adopted in the AGNms model, but it also highlights the need for more theoretical and modeling efforts to accurately reproduce the observed AGN luminosity distribution.
@@
These results offer intriguing insights into the physics of SMBHs and their growth history. Dwarf galaxies typically experience fewer major mergers in their recent assembly compared to their more massive counterparts, making them more sensitive to the “steady mode” of accretion. Our results suggests that a good model for this steady mode could involve accretion tied to the star-formation rate of the host galaxy.
@@ 5 Conclusions
We characterized the sources in our sample by investigating their X-ray hardness ratio, their location within the host galaxy, luminosity function, and galaxy-to-BH mass relation. The average X-ray hardness ratio of our sample is compatible with a power-law model with a slope of $\Gamma=1.8$, as expected from a population of AGN. While the broad PSF of eROSITA prevents us from precisely locating the X-ray sources within their host galaxies, we found that a substantial fraction ($\approx 50\%$) of the sources are likely off-nuclear, in line with theoretical predictions. We also constructed a precise X-ray luminosity function for AGN in dwarf galaxies, which we fit with a power-law function. Finally, we compared the BH mass of our sources (inferred from their X-ray luminosity) with their host galaxy stellar mass. We found that the ratio of BH-to-galaxy mass for our sample is in good agreement with the local relation for AGN hosted in more massive galaxies.
To place our results in context, we compared the mass and luminosity distributions of our sample of AGN in dwarf galaxies with predictions from semi-analytical models. Based on this analysis we concluded that (a) the preferred mode of accretion appears to be one in which the accretion rate is linked to the star-formation rate of the host galaxy; (b) no model can convincingly reproduce the observed distribution of AGN luminosities, suggesting that more modeling effort is needed to better describe the behavior of AGN in dwarf galaxies; and (c) our sample is too shallow to distinguish between different seeding mechanisms, highlighting the necessity of deeper X-ray surveys to convincingly determine the origin of the observed SMBHs.

### v2

# X-Ray Bright Active Galactic Nuclei in Local Dwarf Galaxies: Insights from eROSITA

@@ Abstract
Although supermassive black holes (SMBHs) reside in the heart of virtually every massive galaxy, it remains debated whether dwarf galaxies commonly host SMBHs. Because low-mass galaxies may retain memory of the assembly history of their black holes (BHs), probing the BH occupation fraction of local dwarf galaxies might offer insights into the growth and seeding mechanisms of the first BHs. In this work, we exploit the Western half of the eROSITA all-sky survey (covering $20,000\rm{deg^{2}}$) and compile a catalog of accreting SMBHs in local ($D<200$ Mpc) dwarf galaxies. Cleaning our sample from X-ray background sources, X-ray binaries, and ultraluminous X-ray sources, we identify 74 AGN-dwarf galaxy pairs. Using this large and uniform sample, we derive a luminosity function of dwarf galaxy AGN, fitting it with a power law function and obtaining ${\rm d}N/{\rm d}L_{\textup{X}}=(15.9\pm 2.2)\times L_{\textup{X}}^{-1.63\pm 0.05}$. Measuring the offset between the dwarf galaxies centroids and the X-ray sources, we find that $\approx 50\%$ of the AGN are likely off-nuclear, in agreement with theoretical predictions. We compare the BH-to-stellar mass relation of our sample with the local and high-redshift relations, finding that our sources better adhere to the former, suggesting that local AGN across different mass scales underwent a similar growth history. Finally, we compare our sources with semi-analytical models: while our sample’s shallowness prevents distinguishing between different seeding models, we find that the data favor models which keep SMBH in dwarf galaxies active at a moderate rate, motivating model improvement by comparison to AGN in the dwarf galaxy regime.
@@ 1 Introduction
Alternative methods to identify SMBHs in dwarf galaxies rely on optical, UV, and infrared variability (Baldassare et al., 2018, 2020; Martínez-Palomera et al., 2020; Ward et al., 2022; Arcodia et al., 2024), coronal lines (Satyapal et al., 2007; Cann et al., 2018, 2021; Salehirad et al., 2022; Reefe et al., 2023), radio emission (Mezcua et al., 2019; Reines et al., 2020; Davis et al., 2022), and X-ray observations (Schramm et al., 2013; Lemons et al., 2015; Pardo et al., 2016; Mezcua et al., 2018; Birchall et al., 2020; Latimer et al., 2021; Ohlson et al., 2024; Bykov et al., 2024). These last two methods are particularly effective in identifying AGN due to the high nuclear-to-host contrast that these bands offer: high angular resolution radio observations allow the separation of the AGN from other radio-emitting components of the galaxy, and the X-ray spectrum of AGN is distinct from other stellar X-ray sources (Ho, 2008; Merloni, 2016). However, a sensitive volume-limited radio survey of dwarf galaxies in the nearby universe is prohibitively expensive (Reines et al., 2020). Similarly, X-ray observational campaigns have, to date, been either too shallow and with poor angular resolution, or too limited in sky coverage.
The game-changer in this domain is eROSITA (extended ROentgen Survey with an Imaging Telescope Array, Predehl et al. 2021), the main instrument aboard the German-Russian satellite Spektrum Roentgen Gamma (SRG) (Sunyaev et al., 2021). eROSITA provides today’s most sensitive all-sky survey in the X-ray band, and eRASS1, the first data release, comprises more than 930,000 individual sources across the Western half of the sky (Merloni et al., 2024). In this work, we leverage the opportunity offered by eRASS1 to build a sample of uniformly selected AGN in local dwarf galaxies, thereby probing their properties.
Similar efforts have been conducted by Bykov et al. (2024), who utilized eROSITA data from the Eastern half of the sky and the MPA-JHU galaxy catalog, based on the SDSS DR12 (Alam et al., 2015), and Hoyer et al. (2024), who focused on massive BHs hosted in nuclear star clusters. Based on the same parent samples we adopted, i.e. eRASS1 and the HECATE catalog (Kovlakas et al., 2021), but focused on an entirely different science goal, is the effort performed by Kyritsis et al. (2024). They studied the star-formation-driven X-ray emission of normal galaxies, which we instead excluded from our sample which collects only accretion-powered sources.
@@ 2.3 Unresolved X-ray binaries
The formula provided by Lehmer et al. (2010) derives the luminosity in the $2-10$ keV band, so we converted the $0.2-2.3$ keV luminosity provided by the eRASS1 catalog to the $2-10$ keV band assuming a Galactic absorption of $3\times 10^{20}$ cm-2 and a power law with photon index $\Gamma=2$.
@@ 2.4 Resolved ultra-luminous X-ray sources
Figure 5: A representative subset of our sample of AGN hosted in local dwarf galaxies. The three-color composite images were obtained using the SDSS gri-filters. The X-ray counts are shown with the white (two counts) and red (one count) pixels. The cross and the small magenta circle indicate the centroid and positional uncertainty of the X-ray source taken from the eRASS1 point source catalog (Merloni et al., 2024), respectively (typically around $5\arcsec$). The large cyan circle shows the PSF size ($r=26\arcsec$) of eROSITA in scanning mode.
@@ 2.6 Final sample of dwarf galaxy-X-ray source pairs
Our final sample, after removing the four known ULXs, is composed of 74 dwarf galaxies with an eROSITA-detected source, which corresponds to a sample-average AGN fraction of $74/5775\approx 1.3\%$. In Figure 6 we show the X-ray luminosity of the sources as a function of the redshift of the galaxies. Postcard images of a representative set of three of these candidates are shown in Figure 5. All the relevant properties of each dwarf galaxy-X-ray source pair are reported in Table 2.
@@ 3.2 Off-nuclear SMBHs
These results align with theoretical predictions and observational evidence (Bellovary et al., 2019; Reines et al., 2020; Bellovary et al., 2021) for AGN in dwarf galaxies (as also shown in blue in Figure 7), although some caveats must be considered. The large PSF of eROSITA in scanning mode ($\approx 26\arcsec$) prevents any conclusive statement on possible source blending, and dwarf galaxies often exhibit irregular morphology, making it sometimes difficult to identify their centers (Pasetto et al., 2003; Lazar et al., 2024). To conclusively determine the nature and location of the X-ray sources, follow-up high spatial resolution observations with Chandra would be needed.
@@ 3.4 Galaxy-to-black hole mass relation
Exploiting the analogies between high-redshift AGN and AGN hosted in low-mass galaxies, we further investigate these effects by studying the $M_{\bullet}-M_{\star}$ relation in our dwarf galaxy sample. We infer the SMBH mass from the X-ray luminosity of the AGN. To this end, we employ the bolometric correction $L_{\textup{bol}}/L_{\textup{2-10keV}}=16.7$ (see e.g., Lusso et al. 2012; Duras et al. 2020) and assume that all black holes in our sample accrete at $1\%$ of their Eddington ratio, these choices are commonly used in the literature, however, we point the reader attention to the fact that they are calibrated on larger mass BHs. The inferred black hole masses are in the range of $\sim 10^{4}-10^{7}\ \rm{M_{\odot}}$. Figure 9 shows the $M_{\bullet}-M_{\star}$ relation for the AGN in our sample. The blue line and shaded region indicate the local relation for AGN and its dispersion (Reines & Volonteri, 2015). The red and green lines and shaded regions show the high-redshift relations proposed by Pacucci et al. (2023) and Li et al. (2024), respectively.
Our sample of AGN in dwarf galaxies is compatible with the low-redshift $M_{\bullet}-M_{\star}$ relation, although it shows significant scatter and is biased towards higher luminosities at the low-mass end of the plot (the dotted line indicates the lower limit on the detectable black hole mass). This suggests that local AGN hosted in dwarf galaxies share the same $M_{\bullet}-M_{\star}$ relation as more massive $z=0$ galaxies, in agreement with previous results (Schutte et al., 2019; Ellis et al., 2024).
@@ 4 Comparison with semi-analytical models
Semi-analytical models (SAM) represent a powerful tool for studying BH seeding mechanisms as well as for exploring growth modes of BHs (Somerville et al., 2008; Bower et al., 2010; Ricarte & Natarajan, 2018a, b). In this work, we compare the properties of the sources in our sample with mock catalogs of AGN built following the approach described in Chadayammuri et al. (2023) based on the SAMs developed by Ricarte & Natarajan (2018b).
These SAMs consider two seeding mechanisms and three different mass growth modes, combined to build six mock catalogs. In this model, dark matter merger trees are constructed for a variety of halo masses for $200<z<$ 20. SMBHs are seeded at $z>15$ with either rare “heavy” seeds, typically resulting from direct collapse (DC) scenarios, distributed following Lodato & Natarajan (2006) and Volonteri & Begelman (2010), or abundant “light” seeds with a distribution motivated by cosmological simulations (Hirano et al., 2015). This second scenario is motivated by cosmological simulations where BHs are produced by the death of population III (PopIII) stars and the mass distribution of the seeds is represented by a power law spanning from 30 to 100 solar masses (Hirano et al., 2015). Most growth is merger-driven: major mergers cause SMBHs to grow to the $M_{\bullet}-\sigma$ relation, and a few “steady” modes are explored to prescribe accretion rates at other times, detailed below.
@@
  • Broad Line Quasar (BLQ): When a major merger occurs, central BHs grow at Eddington ratios drawn from observed broad line quasars, allowing occasional super-Eddington accretion. Otherwise, the accretion rate is zero.
Previously, Ricarte & Natarajan (2018b) showed that the PL model produced reasonable AGN luminosity functions for $60<z<$ 6, while the observationally motivated AGNms produced too many faint AGN and too few bright AGN at low redshift. All models reproduce the $M_{\bullet}-\sigma$ relation at high mass by construction. The BLQ model produced even better agreement with AGN luminosity functions (Ricarte et al., 2019). However, Chadayammuri et al. (2023) revealed that this model produced too few faint AGN, due to its lack of a steady mode. Fundamentally, this is due to the fact that luminosity function constraints do not directly test how given AGN luminosities are assigned to individual galaxies. As we shall show, our more detailed study probing luminosity as a function of stellar mass illuminates weaknesses in these models.
@@
We can probe growth modes and seeding mechanisms by comparing our observed AGN properties with the mock catalogs built upon the described SAMs. Following Chadayammuri et al. (2023), for each SAM, we build the corresponding mock catalog by computing the expected X-ray luminosity based on the mass and mass accretion rate of each BH. We assume an efficiency in converting mass accretion rate into luminosity of $10\%$ and a bolometric-to-X-ray luminosity correction of $10\%$, which are both common and adequate choices (see e.g., Lusso et al. 2012; Duras et al. 2020). Although these quantities are poorly understood in dwarf galaxies, these are not the leading sources of theoretical uncertainty. We then filter the obtained mock catalogs using the same luminosity, mass, and distance criteria adopted in the construction of our AGN sample in dwarf galaxies. This approach allows for a direct comparison between the filtered mock catalogs and our observed sample.
Figures 10 and 11 present our main results, showing the normalized distributions of host-galaxy masses and luminosities for our sample (in grey) and the six mock catalogs built on different models (colored dashed lines). Three models roughly reproduce the distribution of observed host-galaxy mass: DC+AGNms, PopIII+AGNms, and PopIII+BLQ. These results suggest that the data prefer models with more abundant AGN in dwarf galaxies. The AGNms model keeps all SMBHs accreting at a moderate rate. While the BLQ model does not have a steady accretion mode, the higher abundance of PopIII seeds keeps the AGN fraction higher than for DC seeds.
On the other hand, the PopIII+BLQ model fails to predict the observed distribution of AGN luminosities, producing AGN that are too bright. Again, this points towards the need for a steady mode of AGN activity independent of galaxy mergers to produce fainter AGN at a greater abundance. The two best-performing models (DC+AGNms and PopIII+AGNms) do produce a wider distribution of luminosities, but they fail to reproduce the observed luminosity distribution. They predict distributions skewed toward higher luminosities, peaking at about $10^{41.5}\ \rm{erg\ s^{-1}}$, while the observed distribution peaks roughly at $10^{39.5}\ \rm{erg\ s^{-1}}$. This suggests that the preferred accretion mode is the one adopted in the AGNms model, but it also highlights the need for more theoretical and modeling efforts to accurately reproduce the observed AGN luminosity distribution. Recall that the AGNms model universally assigns accretion rates at $10^{-3}$ times the SFR, with a duty cycle of unity. Redistributing this accretion in a distribution in future work may help match these observations. This discrepancy is missed when calibrating to the luminosity function of more luminous AGN.
These results offer intriguing insights into the physics of SMBHs and their growth history. Dwarf galaxies typically experience fewer major mergers in their recent assembly compared to their more massive counterparts, making them more sensitive to the “steady mode” of accretion. Our results suggests that a good model for this steady mode could involve accretion tied to the star-formation rate of the host galaxy, or should otherwise keep SMBH active at a moderate rate in dwarf galaxies.
@@ 5 Conclusions
We characterized the sources in our sample by investigating their X-ray hardness ratio, their location within the host galaxy, luminosity function, and galaxy-to-BH mass relation. The average X-ray hardness ratio of our sample is compatible with a power-law model with a slope of $\Gamma=1.8$, as expected from a population of AGN. We found that a substantial fraction ($\approx 50\%$) of the sources are likely off-nuclear, in line with theoretical predictions. We also constructed a precise X-ray luminosity function for AGN in dwarf galaxies, which we fit with a power-law function. Finally, we compared the BH mass of our sources (inferred from their X-ray luminosity) with their host galaxy stellar mass. We found that the ratio of BH-to-galaxy mass for our sample is in good agreement with the local relation for AGN hosted in more massive galaxies.
To place our results in context, we compared the mass and luminosity distributions of our sample of AGN in dwarf galaxies with predictions from semi-analytical models. Based on this analysis we concluded that the preferred models keep SMBH in dwarf galaxies active at a moderate rate. Among the models tested, this corresponds to one in which the accretion rate is kept proportional to the star-formation rate of the host galaxy. However, even these models do not reproduce the observed distribution of AGN luminosities in detail, producing too few AGN with $\log L_{x}\sim 10^{39.5}$ and too many AGN with $\log L_{x}\sim 10^{41}$, motivating additional modeling efforts to better describe the behavior of AGN in dwarf galaxies. These discrepancies can be missed by modelers calibrating only to luminosity functions of more luminous AGN. Unfortunately, due to the degeneracies between seeding and accretion rates and the shallowness of our sample, we cannot convincingly distinguish between different seeding mechanisms. In addition to improved modeling of AGN in dwarf galaxies, this highlights the necessity of deeper X-ray surveys to convincingly determine the origin of the observed SMBHs.
Acknowledgments